Let X have a normal distribution with mean µ = 30 and standard deviation σ = 10. Calculate P(X > 40), and round your answer to two decimal places.
Answer #1

Solution:
Given in the question
X have a normal distribution
Mean = 30
Standard deviation. =10
We need to find
P(X>40) = 1-P(X<40)
Z = (40-30)/10 = 1
From Z table we can found P-value
P(X>40) = 1- 0.8413 = 0.1587 or 15.87%
so there is 15.87% probability tat X is greater than 40.